OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 5? 5 to the power of 1 over 3 5 to the power of 1 over 6 5 to the power of 2 over 3 5 to the power of 3 over 2

OpenStudy (misty1212):

\[\sqrt{\sqrt[3]{5}}\]like that?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

multiply the "indexes" \[2\times 3=6\]

OpenStudy (xapproachesinfinity):

use the rule!

OpenStudy (misty1212):

so \[\sqrt[6]{5}\]

OpenStudy (misty1212):

always use the rule!

OpenStudy (misty1212):

not sure what rule it is then \[\huge \sqrt[6]{5}=5^{\frac{1}{6}}\]

OpenStudy (anonymous):

That is what I was thinking, thanks to both of you!

OpenStudy (anonymous):

Can you guys help me with another one?